OpenStudy (anonymous):

Which equation has solution x = -3? A) 2x - 7 = -1 B) 3x + 8 = 1 C) 1 2 x + 8 = 10 D) 1 2 (2x - 6) = -6

OpenStudy (anonymous):

Do you mind writing this out a little bit better? I am having a hard time reading the question with the random 1's and 2's.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Which equation has solution x = -3? A) 2x - 7 = -1 B) 3x + 8 = 1 C)1/2/ + 8=10 D)1/2(2x-6)=-6

OpenStudy (anonymous):

Oh okay that makes more sense

OpenStudy (anonymous):

Sorry i took so long

OpenStudy (anonymous):

Is C) supposed to be C) 1/2x + 8=10 instead of C) 1/2/ + 8=10

OpenStudy (anonymous):

Yes you are right sorry dear :)

OpenStudy (anonymous):

No biggey.... okay so the easiest way to do this is to just plug in -3 wherever we see x...and if the equation is equal to itself we have the answer... A) 2x - 7 = -1 Multiply 2 adn -3

OpenStudy (anonymous):

A) 2x - 7 = -1 A) 2(-3) - 7 = -1

OpenStudy (anonymous):

Thank you so much , Do you have time for more

OpenStudy (anonymous):

First off, what is the answer?

OpenStudy (anonymous):

Firth A

OpenStudy (anonymous):

First

OpenStudy (anonymous):

No, I am sorry...i think I mislead you... you need to do this for each of the answers! So for the first one, what is 2 times -3?

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

Yes...-6 + -7 is what?

OpenStudy (anonymous):

-13

OpenStudy (anonymous):

Yes, exactly. Which means we now have the equation... -13=-1 -13 IS NOT equal to -1 so A is not the correct answer.

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Now for B... 3x + 8 = 1 What is 3 times -3? (remember this is because wherever we see an x, we plug in -3 for it.

OpenStudy (anonymous):

0

OpenStudy (anonymous):

No, 3 TIMES -3..not 3 + -3

OpenStudy (anonymous):

Oh sorry , -9

OpenStudy (anonymous):

You're cool :) You are also correct -9 + 8 = 1 -9 ADDED to 8 is what? (-9 + 8)

OpenStudy (anonymous):

1

OpenStudy (anonymous):

No, here, do this in a calculator... 8 - 9 and see what you get

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

Yup! -1 = 1 -1 and 1 are not the same thing...so B is not right either

OpenStudy (anonymous):

Eliminated

OpenStudy (anonymous):

C )1/2x + 8 = 10 What is 1/2 TIMES -3/1? Remember that with any number that is NOT a fraction we can divide it by 1 and it will always be the original number. So do 1/2 times -3/1

OpenStudy (anonymous):

Do you understand this?

OpenStudy (anonymous):

Im trying to work it out

OpenStudy (anonymous):



OpenStudy (anonymous):

-3/2

OpenStudy (anonymous):

Yup! C ) -3/2 + 8 = 10 Now comes the part I HATE in math.... So now we need to add a fraction (-3/2) and a normal number (8) So...make 8 become 8/1...but the denominator (the bottom of the fraction) are differnt and we cannot add them this way....so we have to make the denominators the same. Multiply 8/1 by 2/2....so do 8 times 2 and then do 2 * 1

OpenStudy (anonymous):

16/2

OpenStudy (anonymous):

Yes. Now add the numerators (the top of the fractions) together. the denominators will still be 2. What is -3 + 16?

OpenStudy (anonymous):

Or a better way to write this is what is 16 - 3?

OpenStudy (anonymous):

13

OpenStudy (anonymous):

C ) 13/2 = 10 Now divide 13 by two... 

OpenStudy (anonymous):

6r1

OpenStudy (anonymous):

Yup! And 6r1 is NOT EQUAL to 10, so C is also out. That means the only answer can be D.... D) 1/2 (2x-6) = -6 I will do this one for you... D) 1/2 (2(-3)-6) = -6 D) 1/2 (-6 - 6) = -6 D) 1/2 (-12) = -6 D) 1/2 TIMES -12/1 = -6 D) -12/6 = -6 D) -6 = -6 And -6 IS EQUAL TO -6, so D is out answer

OpenStudy (anonymous):

Did you get it